If C=5 and a=3 what does b= and how do you get the result
Answer by Chika Madekwe(1)

Since a^2+b^2=c^2
sustitute the given information
3^2+b^2=5^2
9+b^2=25
Subtract 9 from both sides
b^2=16
Square roots of both sides
b=4